# CodeAssistBench (CAB): Dataset & Benchmarking for Multi-turn Chat-Based Code Assistance

### v2

@@ Abstract
Programming assistants powered by large language models have transformed software development, yet most benchmarks focus narrowly on code generation tasks. Recent efforts like InfiBench and StackEval attempt to address this gap using Stack Overflow data but remain limited to single-turn interactions in isolated contexts, require significant manual curation, and fail to represent complete project environments. We introduce CodeAssistBench (CAB), the first benchmark framework for evaluating multi-turn programming assistance in realistic settings that address real-world questions about actual codebases. Unlike existing programming Q&A benchmarks, CAB automatically generates scalable datasets from question-related GitHub issues using configurable parameters (e.g., repository creation date, star count, programming languages), and includes automatic containerization of codebases for evaluation. It then evaluates models through simulated users in these containerized environments with full codebase access. Using this framework, we constructed a test set of 3,286 real-world programming questions across 231 repositories, spanning seven programming languages and diverse problem domains. Our evaluation of leading LLMs reveals a substantial capability gap: while models perform well on Stack Overflow questions with success rates of 70-83%, they resolve only up to 16.49% of CAB’s recent issues. This discrepancy highlights the challenges of providing assistance in complex, project-specific contexts versus answering standalone questions. Our fully automated framework enables continuous benchmark expansion and is available at https://anonymous.4open.science/r/CAB-CBA3/.
@@ 1 Introduction
Large language models (LLMs) are increasingly used throughout programming workflows. Early benchmarks such as HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) focused on generating functionally correct code snippets for isolated problems. As the field matured, benchmarks like CodeContests Li et al. (2022) introduced harder tasks; while others, including SWE-benchh Jimenez et al. (2023), BigCodeBench Zhuo et al. (2024), and SWE-PolyBench Rashid et al. (2025) began to address more realistic scenarios. Multi-turn benchmarks such as ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) further advanced this direction by incorporating conversational interactions, though they still primarily evaluate code synthesis. Despite these advances, most benchmarks fail to capture the full spectrum of developer needs. A recent Stack Overflow survey of 34,168 developers reveals that when asked about future AI tool expectations, developers expressed strong interest across multiple assistance categories. While 75.7% want AI for code generation, even higher percentages desire AI support for searching for answers (77.9%), debugging and troubleshooting (77.3%), and learning about codebases (73.6%) Stack Overflow (2024). These findings demonstrate that developers anticipate using AI tools as comprehensive collaborators throughout the entire development lifecycle.
Recent benchmarks like InfiBench Li et al. (2024) and StackEval Shah et al. (2024) incorporate Stack Overflow data to better reflect developer Q&A scenarios. However, these benchmarks have critical limitations: they use only single-turn interactions in isolated contexts, require extensive manual curation, and fail to represent real project environments where developers engage in iterative exchanges - trying suggestions, encountering errors, and requesting clarifications. The text clearly states that current language models are achieving high success rates on two benchmarks (InfiBench at 70.64% and StackEval at 83.0%), and shows that these benchmarks no longer represent significant challenges for recent LLMs.
Figure 1: Port conflict clarification in a real-world exchange.
Most multi-turn benchmarks for code generation have been limited to programming tasks, where evaluation typically relies on a binary oracle - running tests to check if they pass. Figure 1, parsed from a real GitHub issue, shows why that criterion fails for programming assistance. Even a trivial port-mapping question forces an assistant to (1) infer project topology, (2) explain that the proxy port is hard-coded, and (3) reassure the user that no extra mapping is needed. Each reply reshapes the next user utterance, so we must judge an evolving conversation, not a single execution. What matters is a gradient of clarity, efficiency, and user confidence that unfolds over turns. Evaluating it demands benchmarks that track satisfaction conditions across the dialogue, not just test outcomes.
To address these limitations, we introduce CodeAssistBench (CAB), a chat-based benchmark framework for evaluating multi-turn programming assistance grounded in real-world developer scenarios. Our contributions include:
1. A fully automated data generation framework that enables continuous benchmark creation and expansion without manual curation. This automation allows us to create dataset variants including recent repositories (post-training cutoff), ensuring the benchmark remains challenging as models evolve.
2. A realistic multi-turn Q&A dataset with 3,286 real-world programming questions across 231 repositories, spanning seven programming languages and diverse problem domains, all derived from actual GitHub issues.
3. A novel chat-driven evaluation pipeline that simulates realistic user-agent interactions in containerized environments with full codebase access. Our pipeline includes a simulated user that provides contextual feedback, a maintainer agent with environment access, and an automated judge that evaluates conversation quality against extracted satisfaction conditions.
Our evaluation of leading LLMs reveals a substantial capability gap: even the best model solve only 16.49% of recent CAB issues compared to 29.14% on historical data. This performance difference highlights both the challenges of CAB and the value of our framework in generating continuously evolving benchmarks. To advance AI coding assistance, we release the complete CAB benchmark, evaluation framework, and detailed results publicly here.
@@ 2 Related Work
Although programming assistance spans diverse domains, current benchmarks fail to reflect the full complexity of real-world developer support.
Code Generation Benchmarks. Traditional evaluations of programming assistants have focused on functional code generation. HumanEval Chen et al. (2021), MBPP Austin et al. (2021), and CodeContests Li et al. (2022) assess model ability to generate functionally correct code from well-specified prompts. More advanced benchmarks such as SWE-Bench Jimenez et al. (2023) and SWE-PolyBench Rashid et al. (2025) use real GitHub issues to evaluate software maintenance tasks across actual repositories. While these provide strong correctness metrics using pass@1, they are largely limited to single-shot interactions and do not capture the iterative nature of real developer workflows.
Multi-turn Programming and Conversational Benchmarks. Several benchmarks aim to evaluate iterative interactions. ConvCodeWorld Han et al. (2025), MINT Wang et al. (2023), and TICODER Fakhoury et al. (2024) simulate developer-assistant dialogues but typically assume consistent environments and remain centered on code synthesis. More general conversational benchmarks—such as MT-Bench Zheng et al. (2023), AlpacaEval Dubois et al. (2023), and BotChat Duan et al. (2024) - offer valuable insights into dialogue quality but do not target the unique demands of technical problem-solving in programming contexts. In contrast, our work directly addresses this intersection by evaluating LLMs in multi-turn, task-grounded developer scenarios across full codebases.
Programming Q&A Benchmarks. Recent benchmarks like StackEval Shah et al. (2024) and InfiBench Li et al. (2024) leverage StackOverflow data to better reflect how developers seek help in practice. StackEval evaluates LLMs on 925 questions across 25 languages and task types, while InfiBench offers 234 examples with model-agnostic evaluation metrics. These benchmarks better capture explanatory reasoning than traditional code benchmarks, but they are still constrained to single-turn formats and require significant manual curation, limiting scalability.
@@ 3 CAB: CodeAssistBench
Figure 2: Overview of CAB’s pipeline for dataset generation, consisting of: (1) Repository Collection from GitHub, filtering by stars, creation date, and license; (2) Filtering of issue conversations using regex and LLM-based criteria; and (3) Structured Data Generation including Build Environment, Satisfaction Conditions, and User Response References. This automated pipeline enables continuous benchmark expansion without manual curation.
CodeAssistBench (CAB) is a comprehensive benchmark framework designed to evaluate large language models (LLMs) in realistic, multi-turn programming assistance scenarios. Unlike prior benchmarks that focus on code generation or single-turn evaluation, CAB simulates developer workflows grounded in GitHub issues. Our framework consists of two key components: (1) an automated dataset generation pipeline that creates diverse, high-quality programming assistance scenarios from GitHub issues; and (2) a multi-agent evaluation framework that simulates realistic user-model interactions in containerized environments.
@@ 3.1 Dataset Generation Pipeline
The CAB dataset generation pipeline automatically transforms GitHub issues into structured programming assistance scenarios through three main stages: repository collection, issue filtering, and data preparation (Figure 2). This section elaborates on the details of this pipeline, while detailed statistics about the output of each phase are presented in Section 4.
@@ 3.1.1 Repository Collection
We begin by collecting a diverse set of GitHub repositories that represent real-world programming challenges across multiple languages. Let ${R}_{GH}$ denote the set of all public GitHub repositories. We extract a filtered subset:
@@
where:
  • $s(r)$ is the star count of repository $r$, with threshold $S_{\min}$ (e.g., 10),
  • $t(r)$ is its creation date, with cutoff $t_{0}$ (e.g., 2024-11-01),
  • $\ell(r)$ is $r$’s SPDX license, restricted to permissive licenses compatible with research use.
To prioritize repositories with active developer support communities, we define a community score:
$\mathrm{CS}(r)=Q(r)+H(r)$ (2)
where $Q(r)$ and $H(r)$ represents the number of closed issues labeled with “question” and “help wanted” respectively. We then select the top-$K$ repositories ranked by $\mathrm{CS}(r)$ (breaking ties by star count) to form our repository set $R_{K}$, where $K$ is a user-defined parameter.
@@ 3.1.3 Data Preparation
For each filtered issue, we prepare three essential components that enable realistic evaluation: build environment generation, user satisfaction condition extraction, and user response reference generation.
[…]
@@
Result Aggregation. The resulting dataset entry for each issue is a tuple:
[…]
For each issue $i\in\bigcup_{r}I_{r}^{\prime}$, where $I_{r}^{\prime}$ denotes the filtered issues for repository $r$, we create such entries to form our complete dataset.
3.2 Evaluation Framework
@@
CAB’s evaluation framework simulates realistic programming assistance interactions through a multi-agent system with three distinct roles: a User, a Maintainer, and a Judge (Figure 3).
User Agent. The user agent initiates the conversation with a programming question extracted from a GitHub issue and provides follow-up responses based on the model’s answers. This agent presents the initial question with relevant context, evaluates model responses against satisfaction conditions, provides realistic follow-up questions or clarifications, and signals when a solution successfully resolves the issue. The user agent does not directly interact with the environment but observes execution results to assess progress toward resolution.
Maintainer Agent. The maintainer agent represents the LLM being evaluated. Given the user’s question and contextual information, the model must analyze the problem within the provided codebase context, execute commands in the containerized environment when necessary, generate helpful responses, and adapt its approach based on user feedback. To balance exploration with efficiency, the maintainer can take up to 5 exploration steps per response, where each step may involve file operations or command execution.
Judge Agent. After the conversation concludes, an automated judge evaluates the interaction quality based on: (1) technical correctness - whether the proposed solution correctly addresses the underlying issue; (2) satisfaction completeness - whether all extracted satisfaction conditions are met; and (3) interaction quality - whether the conversation was appropriately concise and helpful. For issues with Docker environments, execution success is a hard requirement - a technically sound explanation is insufficient if the implementation fails in practice. The evaluation is triggered when either the user expresses satisfaction or the conversation reaches a maximum of 10 turns. This approach provides a comprehensive assessment of an LLM’s ability to provide effective programming assistance in realistic scenarios.
3.3 Implementation Details
CAB is implemented as a fully automated pipeline that can continuously generate new benchmark datasets as GitHub repositories evolve. Our implementation integrates with GitHub’s REST API to collect repositories and issues with configurable parameters, automatically generates and validates Docker environments to ensure reproducibility, and orchestrates the multi-agent conversation flow while recording detailed interaction logs. This implementation enables CAB to generate continuously evolving benchmarks that remain challenging as models improve, particularly through the creation of dataset variants that include repositories created after model training cutoff dates.
@@ 4 Experimental Results and Analysis
In this section, we report statistics from our dataset generation pipeline, model evaluation experiments, and human alignment analysis. For details on the hardware setup and compute time associated with these experiments, we refer readers to Appendix D.
@@ 4.1 Statistics of Dataset Generation Pipeline
We applied our dataset generation pipeline (Section 3.1) to two distinct repository cohorts - All-Time and Recent - to evaluate its scalability and effectiveness. The All-Time cohort comprises 700 top-starred repositories (100 per language) without creation date constraints, from which we selected 70 (10 per language) based on permissive licenses and community engagement. The Recent cohort includes 3,500 top starred repositories (500 per language) created after November 2024, from which 700 were selected using the same criteria - yielding a total pool of 770 repositories across seven programming languages  [^1] There is no overlap between the All-Time and Recent cohorts.
From these repositories, CAB collected 25,656 raw GitHub issues with question and help-wanted issue tags. A combination of regex - and LLM-based filters was used to discard issues with insufficient content or ambiguous resolutions. A summary of the filtered issues, along with Docker requirement classification, build results, and final issue retention, is presented in Table 1. Recent repositories exhibited a notably higher Docker build success rate - 97.6% (41 out of 42) - compared to 78.2% (197 out of 252) in All-Time repositories. Most build failures in older repositories stemmed from outdated dependencies or unavailable components, not from limitations of the LLM pipeline itself. To address this, CAB inferred project-specific requirements - such as the relevant library versions, operating systems, and toolchains - by analyzing both the issue’s content and its creation timestamp, and then tailored environment reconstruction accordingly via custom Dockerfile generation.
Ultimately, our dataset includes 3,286 multi-turn conversations from 214 repositories. Among these, 294 issues required Docker-based environments, and 238 were successfully built through our automated pipeline. A total of 44,628 Sonnet 3.7 calls were made across all steps, without any manual intervention after the initial setup. Language-wise, recent repositories skew toward Python and TypeScript, while All-Time repositories present a more balanced distribution.
As shown in Figure 5, over half of the issues across all languages involved multiple conversational turns. In particular, Python, C++, and TypeScript showed the highest rates of extended discussions (3+ turns), underscoring the need for multi-turn understanding and evaluation in real-world software development scenarios.

### v3

@@ Abstract
Programming assistants powered by large language models have improved dramatically, yet existing benchmarks still evaluate them in narrow code-generation settings. Recent efforts such as InfiBench and StackEval rely on Stack Overflow questions and remain limited to single-turn interactions, manually curated data, and isolated snippets rather than full project environments. We introduce CodeAssistBench (CAB), the first benchmark for evaluating multi-turn, project-grounded programming assistance at scale. CAB automatically constructs datasets from GitHub issues tagged as questions, using an LLM-driven pipeline that filters noise, extracts runnable contexts, builds executable containers, and verifies environment correctness. This enables continuous, automated expansion across diverse repositories without manual intervention. Using CAB, we create a testbed of 3,286 real-world issues across 214 repositories, spanning seven languages. Evaluating state-of-the-art models reveals a substantial gap: while models achieve 70–83% accuracy on Stack Overflow–style questions, they solve only 16.49% of CAB issues from post-training-cutoff repositories. On a manually validated subset of 149 issues, top models such as Claude Sonnet 4.5 reach only 12.08% correctness. These results highlight a fundamental challenge: current LLMs struggle to provide assistance in realistic, project-specific contexts despite strong performance on traditional Q&A benchmarks. CAB provides a scalable, reproducible framework for advancing research in multi-turn, codebase-grounded programming agents. The benchmark and pipeline are fully automated and publicly available at https://github.com/amazon-science/CodeAssistBench/.
@@ 1 Introduction
Large language models (LLMs) are increasingly integrated into modern programming workflows, supporting tasks ranging from code generation to debugging and repository navigation. While benchmarks have progressed from isolated synthesis tasks (HumanEval chen2021evaluating, MBPP austin2021program) to repository-level maintenance settings (SWE-Bench jimenez2023swe, BigCodeBench zhuo2024bigcodebench), they still capture only a narrow slice of how developers seek assistance in practice. Recent multi-turn benchmarks such as ConvCodeWorld han2025convcodeworld, MINT wang2023mint, and TICODER fakhoury2024ticoder move toward conversational evaluation, yet they primarily emphasize code synthesis and assume stable, well-defined contexts.
However, real-world programming assistance extends far beyond writing code. A 2024 Stack Overflow survey of 34,168 developers reports that while 75.7% seek AI for code generation, even higher percentages desire support for searching for answers (77.9%), debugging and troubleshooting (77.3%), and understanding unfamiliar codebases (73.6%) stackoverflow2024aisurvey. These tasks require iterative clarification, environment-aware reasoning, and integration with project-specific details—capabilities that current benchmarks largely overlook.
Recent efforts such as InfiBench li2024infibench and StackEval shah2024stackeval incorporate Stack Overflow data to better reflect developer Q&A scenarios, but they remain limited to single-turn interactions over isolated snippets and require substantial manual curation. Critically, they do not model how developers and assistants interact over multiple turns: trying suggestions, encountering new errors, revising assumptions, and progressively forming shared understanding.
Figure 1: A port-mapping clarification from a real GitHub issue.
Real assistance is inherently interactive. Even seemingly simple questions often require the assistant to infer project topology, reference environment configuration, or reconcile ambiguous user descriptions. Figure 1 illustrates a real GitHub issue in which a user asks about container port mapping. Solving it requires the assistant to (1) interpret the repository’s network structure, (2) explain that the proxy port is internal and hard-coded, and (3) assure the user that no additional mapping is needed. Each answer influences the next question, defining a trajectory of reasoning and communication. Evaluating such scenarios cannot rely solely on execution oracles or single-shot correctness; it requires tracking how well the assistant guides the user across turns with clarity, efficiency, and accuracy.
To address these limitations, we introduce CodeAssistBench (CAB), the first benchmark designed to evaluate multi-turn, project-grounded programming assistance at scale. CAB automatically constructs datasets from GitHub issues tagged as questions using an LLM-driven pipeline that filters noisy issues, extracts runnable contexts, builds containerized environments, and verifies their correctness. This automation enables continuous, reproducible benchmark expansion without manual curation and supports dataset variants drawn from repositories created after major model training cutoffs, ensuring persistent difficulty over time.
CAB consists of 3,286 real-world programming questions across 214 repositories in seven languages, paired with an execution-backed, chat-driven evaluation pipeline. A simulated user provides contextual feedback, a maintainer agent interacts with the full codebase in a containerized environment, and an automated judge assesses conversation quality using extracted satisfaction conditions. This framework captures the full complexity of real-world developer–assistant interactions: ambiguous problem descriptions, evolving hypotheses, and multi-turn negotiation of understanding.
Our evaluation of leading LLMs reveals a striking capability gap. While recent models achieve 70–83% accuracy on Stack Overflow-style questions, they solve only 16.49% of CAB issues from post-training-cutoff repositories. Even on a manually validated subset of 149 issues, top models reach only 12.08% correctness. These results highlight a fundamental challenge: despite impressive performance on traditional Q&A benchmarks, LLMs struggle to provide reliable assistance in realistic, multi-turn, project-specific contexts.
CAB provides a scalable and reproducible foundation for advancing research in multi-turn programming assistance. By coupling fully automated dataset generation with environment-grounded conversational evaluation, it offers a new lens through which to measure and improve the real-world capabilities of LLM-based developer tools.
@@ 2 Related Work
Although programming assistance spans diverse developer needs, existing benchmarks only capture narrow aspects of real-world workflows.
Code Generation Benchmarks. Traditional evaluations of programming assistants focus on generating functionally correct code from well-specified prompts. HumanEval chen2021evaluating, MBPP austin2021program, and CodeContests li2022competition assess single-shot code synthesis, while more advanced benchmarks such as SWE-Bench jimenez2023swe and SWE-PolyBench rashid2025swepolybenchmultilanguagebenchmarkrepository use real GitHub issues to evaluate repository-level maintenance tasks. These benchmarks provide robust execution-based correctness metrics but remain predominantly single-turn and do not capture the iterative, conversational nature of real developer interactions.
Multi-turn Programming and Conversational Benchmarks. Several benchmarks incorporate iterative interactions. ConvCodeWorld han2025convcodeworld, MINT wang2023mint, and TICODER fakhoury2024ticoder simulate developer–assistant dialogues but typically assume stable, fully specified environments and remain centered on code synthesis tasks. More general conversational evaluations—such as MT-Bench zheng2023judging, AlpacaEval dubois2023alpacafarm, and BotChat duan-etal-2024-botchat—measure dialogue quality but are not targeted at technical problem-solving in programming contexts. In contrast, we evaluate multi-turn, task-grounded interactions where assistance depends on project state, ambiguous context, and evolving user needs.
Programming Q&A Benchmarks. StackEval shah2024stackeval and InfiBench li2024infibench use Stack Overflow data to better reflect practical developer questions and emphasize explanatory reasoning. However, they remain limited to single-turn settings, require substantial manual curation, and lack integration with executable project environments. These constraints prevent them from capturing the iterative reasoning, context reconstruction, and environment-aware troubleshooting that dominate real-world programming assistance.
Overall, prior work evaluates code generation, multi-turn dialogue, or Q&A reasoning in isolation. CAB is the first benchmark to combine multi-turn conversational evaluation with full-project, environment-grounded contexts generated entirely automatically from real GitHub workflows.
@@ 3 CAB: CodeAssistBench
Figure 2: CAB’s automated data generation pipeline. It collects relevant GitHub repositories, filters issue conversations, and produces structured assistance scenarios with build environments, satisfaction conditions, and reference user responses.
CodeAssistBench (CAB) is a fully automated benchmark framework for evaluating multi-turn, project-grounded programming assistance. CAB converts real GitHub issues into structured, executable assistance scenarios and evaluates LLMs through simulated multi-agent interactions. The framework has two components: (1) an automated dataset construction pipeline and (2) an environment-grounded conversational evaluation system.
@@ 3.1 Dataset Generation Pipeline
CAB transforms raw GitHub issues into structured assistance tasks through three stages: (1) repository collection, (2) issue filtering, and (3) data preparation. An overview is shown in Figure 2.
@@ 3.1.1 Repository Collection
We begin by collecting a diverse set of GitHub repositories that represent real-world programming challenges across multiple languages. Let $\mathcal{R}_{GH}$ denote the set of all public GitHub repositories. We extract a filtered subset:
@@
where $s(r)$ is the star count of repository $r$ with threshold $S_{\min}$ (e.g., 10), $t(r)$ is its creation date with cutoff $t_{0}$ (e.g., 2024-11-01), and $\ell(r)$ is $r$’s SPDX license, restricted to permissive licenses compatible with research use.
To prioritize repositories with active developer support communities, we define a community score $\mathrm{CS}(r)=Q(r)+H(r)$, where $Q(r)$ and $H(r)$ represents the number of closed issues labeled with “question” and “help wanted” respectively. We then select the top-$K$ repositories ranked by $\mathrm{CS}(r)$ (breaking ties by star count) to form our repository set $R_{K}$, where $K$ is a user-defined parameter.
@@ 3.1.3 Data Preparation
For each filtered issue, we prepare three essential components: build environment generation, user satisfaction condition extraction, and user response reference generation.
Build Environment Generation. For issues requiring environment-specific testing, we automatically generate Docker configurations by analyzing repository artifacts using Sonnet 3.7 claude37sonnet with a structured prompt (Appendix A.3) that identifies the commit $\mathit{sha}_{i}$ closest to the issue creation timestamp, clones the repository at that commit, and extracts key artifacts (README content, Dockerfiles, GitHub workflows, file structure). It then generates and tests candidate build scripts until finding a successful configuration $e_{i}$.
@@
Result Aggregation. The resulting dataset entry for each issue is a tuple $d_{i}=\bigl(r_{i},\;i,\;\mathit{sha}_{i},\;e_{i},\;s_{i},\;u_{i}\bigr)$. For each issue $i\in\bigcup_{r}I_{r}^{\prime}$, where $I_{r}^{\prime}$ denotes the filtered issues for repository $r$, we create such entries to form our complete dataset.
3.2 Evaluation Framework and Implementation
@@
CAB’s evaluation framework simulates realistic programming assistance interactions through a multi-agent system with three distinct roles: a User, a Maintainer, and a Judge (see Appendix Figure 3).
User Agent. The user agent initiates the conversation with a programming question from a GitHub issue and provides follow-up responses. It presents the initial question with context, evaluates model responses against satisfaction conditions, provides realistic follow-ups or clarifications, and signals when a solution resolves the issue. The user agent observes execution results but does not directly interact with the environment.
Maintainer Agent. The maintainer agent represents the LLM being evaluated. Given the user’s question and contextual information, the model must analyze the problem within the provided codebase context, execute commands in the containerized environment when necessary, generate helpful responses, and adapt its approach based on user feedback. To balance exploration with efficiency, the maintainer can take up to 5 exploration steps per response (configurable), where each step may involve file operations or command execution.
Judge Agent. After the conversation concludes, an automated judge evaluates the interaction quality based on: (1) technical correctness - whether the proposed solution correctly addresses the underlying issue; (2) satisfaction completeness - whether all extracted satisfaction conditions are met; and (3) interaction quality - whether the conversation was appropriately concise and helpful. For issues with Docker environments, execution success is a hard requirement - a technically sound explanation is insufficient if the implementation fails in practice. The evaluation is triggered when either the user expresses satisfaction or the conversation reaches a maximum of 10 turns (configurable). This approach provides a comprehensive assessment of an LLM’s ability to provide effective programming assistance in realistic scenarios.
CAB is implemented as a fully automated pipeline that continuously generates new benchmark datasets as GitHub repositories evolve. Our implementation integrates with GitHub’s REST API to collect repositories and issues with configurable parameters, automatically generates and validates Docker environments to ensure reproducibility, and orchestrates the multi-agent conversation flow while recording detailed interaction logs. This enables CAB to generate continuously evolving benchmarks that remain challenging as models improve, particularly through dataset variants including repositories created after model training cutoff dates.
@@ 4 Experimental Results and Analysis
We evaluate CAB along three dimensions: (1) dataset generation outcomes, (2) model performance in multi-turn debugging, and (3) human studies validating judge reliability and satisfaction-condition quality. Hardware details are provided in Appendix E.
@@ 4.1 Statistics of Dataset Generation Pipeline
We applied our automated pipeline (Section 3.1) to two repository cohorts: an All-Time set of 700 top-starred repositories (100 per language) without creation-date limits, and a Recent set of 3,500 repositories (500 per language) created after November 2024. After applying license and community filters, we retain 770 repositories for downstream processing. These produced 25,656 raw GitHub issues tagged with question or help-wanted. We apply a layered filtering process, regex heuristics followed by an LLM classifier, to remove low-signal or ambiguous issues.
Across both cohorts, we retained 3,286 multi-turn issues from 214 repositories, including 238 fully validated Docker environments. Notably, Recent repositories achieved a 97.6% build success rate compared to 78.2% for All-Time repositories (details in Appendix F). In total, our pipeline executed 44,628 Sonnet 3.7 calls end-to-end without any manual intervention. Issue-length distributions (Figure 5) show that over half of all issues involve multi-turn threads, particularly in Python, C++, and TypeScript—highlighting the need for multi-step reasoning benchmarks such as CAB.